A particle moving along positive x-axis with a speed of 5m/s suddenly changes its direction along the positive y-axis with the same speed.The change in velocity of the particle is what?        (In these type of qstns i am unable to determine the resultant direction of the particle.Can u suggest a simple way of determining it.)

Grade:Upto college level

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Chandan das

here you should see what is initial and final velocity.

initial velocity = 5i

final velcity  = 5j

 

so change in velocity = (final -inital)

                               =5j-5i

   and for magitude = |5j-5i|

                             =5√2

resultant direction is changes with respect to time.

 suppose for t1 sec it moves in x direction and then for t2 sec it moves in the y direction so

x=5t1

y=5t2

so resultant direction = tanΘ=y/x = 5t2/5t1

                                                =t2/t1

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin